OpenStudy (anonymous):

A standard license plate in California has a one digit number, 3 letters, and a 3-digit number, in that order. How many possible standard license plates are there in California

OpenStudy (anonymous):

\[10\times 26\times 26\times 26\times 10\times 10\times 10\] better known as \[26^3\times 10^4\] by the counting principle